Find the measure of the indicated exterior angle.

angle x= (6n - 20)°, angle y = (4n + 47)°, angle z = (170 - n)°
The answers to chose from are:
A) 55°
B) 157°
C) 167°
D) 145°


Answer by nyc_function(2741)

The exterior angle theorem is a theorem in elementary geometry which states that the exterior angle of a triangle is equal to the sum of the two remote interior angles.

Since angle z is the exterior angle, we have the folowing equation: z = x + y.

170 - n = 6n - 20 + 4n + 47

170 - n = 10n + 27

170 - 27 = 10n + n

143 = 11n

143/11 = n

13 = n

z = 170 - n

z = 170 - 13

z = 157

The answer is choice B.